### v1

# Learning in Markets with Heterogeneous Agents: Dynamics and Survival of Bayesian vs. No-Regret Learners

Abstract
We provide an analysis of the performance of heterogeneous learning agents in asset markets with stochastic payoffs. Our agents aim to maximize the expected growth rate of their wealth, but have different theories on how to learn to do this best. Our main focus is on comparing Bayesian learners and no-regret learners in market dynamics. Bayesian learners with a prior over a finite set of models that assign positive prior probability to the correct model have posterior predicted probabilities that converge exponentially fast to the correct model. Consequently, such Bayesians survive even in the presence of agents who invest according to a correct model of the stochastic process. Bayesian learners with a continuum prior converge to the correct model at a rate of $O((\log T)/T)$.
Online learning theory provides no-regret algorithms for maximizing the log of wealth in this asset market setting, achieving a worst-case regret bound of $O(\log T)$ without assuming that there is a steady underlying stochastic process, but comparing to the best fixed investment rule. This regret, as we observe, is of the same order of magnitude as that of a Bayesian learner with a continuum prior. However, we show that even such low regret may not be sufficient for survival in asset markets: an agent can have regret as low as $O(\log T)$, converging to the correct model at a rate of $O((\log T)/T)$, but still vanish in market dynamics when competing against an agent who always invests according to the correct model, or even against a perfect Bayesian with a finite prior. On the other hand, we show that Bayesian learning is fragile, while no-regret learning requires less knowledge of the environment and is therefore more robust. Specifically, any no-regret learner will drive out of the market an imperfect Bayesian whose finite prior or update rule has even small errors. In our analysis, we formally establish the relationship between the notions of survival, vanishing, and market domination studied in economics and the framework of regret minimization, thus bridging these theories. More broadly, our work contributes to the study of dynamics with heterogeneous types of learning agents and their impact on markets. Our results highlight the importance of exploring this area further.
1 Introduction
Learning how to invest in asset markets has been a central topic of study in both the economics and finance literature, as well as in computer science, where algorithmic approaches to decision-making and market dynamics have received growing attention. Generally, players in markets aim to maximize the utility they gain from their resulting wealth. A standard utility function used in both economics and computer science is the $\log$ of the player’s wealth, leading to the objective of maximizing the growth rate of wealth [^1] The economics literature on market selection also considers different objectives based on expected discounted utility of consumption. We consider asset markets in isolation and thus focus on growth rate. See the discussion in Sections 1.1 and 4. (see references below).
However, the two literatures take different approaches to learning—how learning agents evaluate the feedback they receive and what methods they use to select their strategies—and they ask very different questions about the market dynamics. The literature in economics primarily focuses on Bayesian learning, while the computer science literature focuses on the framework of no-regret learning. While both Bayesian and no-regret learners modify their investment decisions over time, the assumptions about the learners and the measures of success are different.
The economics literature studies Bayesian learning, assuming that agents have a prior about the underlying process and update their beliefs over time. If these agents correctly solve their decision problems, they succeed according to their own objectives. In the case that the agents’ objective is expected growth rate and they reach correct beliefs, this aligns also with the external measure of success based on their wealth share in the market. Notably, however, in this approach learners do not directly respond to changes that they experience in their wealth, but rather learn about the underlying stochastic process and best respond to it according to their current beliefs, combining their prior with their experience.
The computer science literature makes no assumptions about the underlying process, allowing it to be adversarial, and focuses on no-regret learning, measuring the resulting outcome by the regret in growth rate compared to the best single strategy in hindsight; that is, the constant investment rule an investor would have used had they known the realization of the market variables. In the case of a market with a steady underlying process that we consider, this includes a comparison to the investor investing knowing the process (although this may not be the best strategy in hindsight for a given history; see Sections 3 and 5). Learning agents in this literature explicitly respond to the rate of change in their wealth at each step, considering it as the utility (or loss) input to their no-regret online learning algorithm.
We study repeated interactions in asset markets with heterogeneous learning agents: both Bayesians and regret minimizers are present in the market and aim to maximize their expected growth rate of wealth. A question that naturally arises in this setting is whether one type of agent drives the other type out of the market. That is, does an agent’s wealth share converge to zero (the agent vanishes) or does it remain bounded away from zero (the agent survives)? A version of this question has been studied in the economics literature in the context of the market selection hypothesis. The results there show that agents who do not correctly maximize expected growth rate vanish in the presence of agents who do correctly maximize expected growth rate. (see the related work section for further discussion and references to this literature).
We ask the market selection question for a market populated with a greater variety of learning agents. Some are Bayesian optimizers, some are Bayesians who make small mistakes in the set of models they consider or in updating their beliefs, and some use no-regret learning. A key question that arises is: what is the meaning or role of regret in this context? Regret-minimizing agents achieve low regret in the complete asset market we analyze, but we show that if there is a Bayesian who makes correct updates and whose prior on models puts positive probability on the correct model, then regret minimizers vanish. That is, surprisingly, successfully reaching low regret does not guarantee survival. However, we also show that Bayesian learning is fragile; if the Bayesian does not have the correct model within the support of its prior, or if the Bayesian does not correctly balance the prior and likelihood functions in updating, then the Bayesian almost surely vanishes relative to the no-regret leaner, even if the Bayesian makes arbitrarily small errors.
1.1 Related Work
No-regret learning: No-regret learning is a fundamental concept in the study of learning in games, tracing back to early work in game theory [14, 24, 49, 71] and subsequently developed in influential papers on no-regret algorithms [8, 16, 42, 44, 50, 59]. See [26, 51, 75] for broad introductions. The dynamics of no-regret learning algorithms in repeated games have been extensively studied in various domains, including repeated auctions [1, 12, 34, 46, 39, 35, 62, 64], budgeted auctions [10, 38, 40, 41, 66], portfolio selection (as we discuss below) [15, 45, 55, 65], repeated contracting [33, 47, 78], bilateral trade [28, 27], and implications of no-regret algorithms on user incentives in general games [61]. In all these works on learning dynamics, either a single learner is considered, or there are multiple learners where it is assumed that all are regret minimizers.
When optimizing growth rate of wealth over multiple interactions, the resulting growth rate (the change in the $\log$ of wealth) is an additive function over the periods, so classical no-regret learning applies. In the context of investment portfolio selection, [15, 55] introduce algorithms with worst-case regret bounds of $O(\log(T))$. See also [26], Chapter 10 and [56] Chapter 4 for a discussion and analysis of investment strategies achieving this regret bound. These studies emphasize regret as the key metric, and in that, implicitly assume that if the time-average of the regret vanishes, then the agent is considered successful. However, we show that in a competitive market, an agent may achieve vanishing regret yet still be driven out of the market, with wealth share converging to zero.
Bayesian learning: In economics, it is standard to view learning as a by-product of expected utility maximization. That is, the decision-making agent’s objective is not learning per se but making good decisions as evaluated by expected utility. It is well known, though not often written formally, that expected utility theory, and its subjective axiomatic foundation, as in [74], imply Bayesian learning. So the focus on Bayes is not surprising.
There is also a long tradition in economics and finance of arguing that the reason financial markets price assets correctly is that the forces of market selection drive out irrational traders, leaving rational traders to determine prices. This argument is historically most closely associated with [2], [43] and [36]. Here, rationality is interpreted as having a statistically correct model of the stochastic process that determines the returns on assets and optimizing according to it. [^2] No-regret learning, in comparison, can be viewed from this perspective as a form of bounded rationality. As this notion of rationality implies Bayes, the argument is often extended to argue that markets select for Bayesians and against all other learning procedures that do not at least asymptotically behave like Bayes. The intuition for the ‘market selection’ argument is simple: if rational and irrational traders disagree, then rational traders take money away from irrational ones in the events that rational traders consider to be likely. The argument then goes on to note that as rational traders compute conditional probabilities correctly, their view of likely events is at least asymptotically correct. So they survive and irrational traders do not survive. This argument has several important qualifications, most importantly, complete markets and the possibility of eventually-correct beliefs for rational traders. The market selection hypothesis, and qualifications to it, have more recently been investigated in [17], [73] and [18].
There is also a literature in finance and operations research on wealth accumulation for investors who maximize the expected growth rate of their wealth. The investment rule that implements this objective is known as the Kelly rule, see [60]. There is a long standing controversy on the applicability of this rule in consumption-investment decisions, see [72] or [57] for a treatment of wealth dynamics with a variety of objectives including expected growth rate maximization.
Interestingly, the literature in computer science on no-regret learning and the literature in economics on Bayesian learning in asset markets differ not only in the questions they ask and methods they use, but are also almost entirely disjoint, with very few cross-references between them. In this work, we establish a formal connection between these different approaches to learning in markets. We show that there is a surprisingly simple relationship between the entropy of the learner—a key notion in the Bayesian learning literature—and the regret of the learner in these markets. We then analyze a heterogeneous group of Bayesian learners and regret minimizers interacting in a competitive market, and study the implications of these different learning approaches for survival in the market.
Dynamics in markets and investment games: A rich body of literature explores the dynamics of various market models. Prominent examples include studies on the tâtonnement price dynamics [29, 32, 76], proportional response dynamics in Fisher markets [30, 63, 77], dynamics in exchange economies [19, 21] and the von Neumann model of expanding economies [20, 22], repeated Cournot competition games [13, 37, 61], and dynamics in pricing games [11, 52, 53]. These studies focus on the joint dynamics of agents, analyzing their convergence properties, equilibria, and implications for overall outcomes. While many of these dynamics are (or can be framed as) distributed learning processes, these works assume homogeneous learning behavior among agents and do not address settings where learners use fundamentally different learning approaches. Other research on investment games focuses on equilibrium outcomes and interventions designed to incentivize investment [9, 48, 58], which do not involve learning.
Games with heterogeneous learners: The present work studies interactions between heterogeneous learning agents and their implications for market dynamics and outcomes. The literature on learning in games contains very few studies that analyze the dynamics between learners using different strategies. The most closely related work to ours is [17], which show that Bayesian learners outperform dynamic heuristics based on imitation or search in investment games. An additional line of work on repeated games where two players use different types of reasoning is on Stackelberg games where an optimizer seeks the best strategies against an opponent who commits to using a fixed no-regret algorithm in the game [5, 6, 7, 23, 25, 47, 67]. This differs significantly from our study, where we do not assume that any player is strategically optimizing as a Stackelberg leader.
1.2 Roadmap and Summary of Our Results
To the best of our knowledge, this work is the first to analyze the dynamics of no-regret learners against agents using other theories of learning, to study the survival of no-regret learners in investment scenarios (in contrast to only their regret level), and to establish a connection between regret minimization and Bayesian learning in investment settings.
We show that, on the one hand, a Bayesian learner dominates the market against any player whose regret increases with time (including logarithmic regret). Such a player may have vanishing regret, but will also have a vanishing share of the market value. On the other hand, we show that Bayesian learning is fragile. If the Bayesian learner is not perfect but, instead, has some level of noise in its prior or update rule, it may initially gain a large share of the market, but in the long run, a no-regret learner will dominate, driving the Bayesian out of the market.
The paper is structured as follows. In Section 2 we present the market model and provide useful preliminary analysis, discussing relative wealths and the role of relative entropy in characterizing survival for stationary investment rules. Section 3 analyzes the relation between regret and wealth shares in the market, showing how regret can be used to compare different learners. In Section 4, we restate in the language of our market setting several known analyses of Bayesian learning that will be important for our comparison of Bayesian learners, no-regret learners, and imperfect Bayesians. In Section 5, we analyze the regret of a player who knows the correct stochastic model of market states and show that this analysis leads to a characterization of survival conditions in competition with a perfect Bayesian. Section 6 studies imperfect Bayesians, showing that Bayesian learners with inaccurate priors or updates vanish when competing against no-regret learners that converge to the correct model. Finally, in Section 7, we present and discuss simulation results for several scenarios of competition between no-regret learners, perfect Bayesians, and imperfect ones, to complement our analysis and provide further intuition.
We conclude the introduction with an a informal summary of our main theorems and analyses:
Regret and Survival: Theorem 3.1 characterizes the regret condition for survival in a competitive asset market: an agent survives if and only if its regret remains bounded by a constant relative to every competitor at all times. Remarkably, if two players both have similar regret asymptotics with respect to the best strategy in hindsight (e.g., both have $O(\log T)$ regret) but with different coefficients, the agent with the larger constant almost surely vanishes from the market.
Regret of the Optimal Investment Strategy: Regret with respect to the best strategy in hindsight is a strong benchmark; as we discuss in Section 3, due to the stochasticity in market dynamics, even the best implementable strategy (i.e., one that does not depend on future outcomes) has positive regret. In Theorem 5.2, in Section 5, we calculate this expected regret level, showing that it is constant and depends only on the number of assets in the market. As a corollary of this result and Theorem 3.1, any agent who survives in competition with an agent who invests knowing the true distribution of states must have regret bounded by a constant at all times.
Survival against Perfect Bayesians: Theorem 5.4 shows that perfect Bayesian learners with a finite support have constant expected regret. Consequently, a learning agent survives in competition with perfect Bayesians if and only if it also has constant regret. The proof uses Proposition 4.2 about Bayesian learners together with the results described above.
Bayesians with Inaccurate Priors: Theorem 6.1 shows that a Bayesian agent with the correct state distribution not within its support suffers linear regret, and, as a result, vanishes from the market in competition with any no-regret learner. This theorem is followed by an analysis of the expected survival times of Bayesians with inaccurate priors, as a function of the size of the errors in the models considered in their priors and the regret rate of the competitor.
Bayesians with Inaccurate Updates: In Section 6.2, we analyze a scenario where a Bayesian learner performs “trembling hand” updates, making small zero-mean errors in the weight given to the current prior and the new data at each step. Theorem 6.5 demonstrates that even such tiny errors break Bayesian learning, causing the learner to fail to converge. As a result, any no-regret (or other) learner that converges to the correct model at any rate outperforms such an inaccurate Bayesian, driving them out of the market.
2 Model and Preliminaries
We use the market structure as in [17] in our analysis of wealth and price dynamics. At each time in discrete steps indexed by $t$, a state is realized. There is a finite number of states $s\in\{1,\dots,S\}$ distributed i.i.d. with probability $q=(q_{1},\dots,q_{S})$ where $q_{s}\geq\Delta>0$ for all $s$. At each time, there is a complete set of assets. Assets are also indexed by $s$ with one unit of asset $s$ purchased at time $t$ paying one unit of wealth at time $t+1$ if state $s$ occurs at time $t+1$ and 00 otherwise. These assets, known as Arrow securities, are more general than they may seem; any asset structure that permits arbitrary reallocation of wealth across states can be written as a simple transformation of Arrow securities. This asset market model, originally introduced in [4], is the standard general equilibrium model of an economy with a complete set of assets, see [3] Chapter 19. Asset prices at time $t$ are represented by $p_{t}=(p_{1t},\dots,p_{St})$.
There are $N$ agents indexed by $n=1,\dots,N$. The wealth of agent $n$ at time $t$ is $w^{n}_{t}$. Agents begin with initial wealths $w^{n}_{0}>0$ for agent $n$. These initial wealths are normalized to sum to one. [^3] This normalization is done only for convenience. We analyze relative wealths. The actual size of the economy (the aggregate wealth) does not play a role in relative wealth dynamics and may follow an arbitrary stochastic process, as long as it is bounded away from zero and bounded from above. Thus, if state $s$ occurs at time $t+1$, agent $n$ will have time $t+1$ wealth, $w^{n}_{t+1}=\frac{\alpha^{n}_{st}w^{n}_{t}}{p_{st}}$.
In a market equilibrium, the price of asset $s$ must be such that the aggregate demand for asset $s$ equals the aggregate supply of one. So the price is given by
$p_{st}=\sum_{n}\alpha^{n}_{st}w^{n}_{t}.$ (1)
Note that asset prices sum to one as the aggregate wealth is one.
Using this structure, we can describe the evolution of wealths and asset prices as functions of the agents’ investments. The step-by-step stochastic process of agent $n$’s wealth is
$w^{n}_{t+1}=\prod_{s=1}^{S}\bigg{(}\frac{\alpha^{n}_{st}}{p_{st}}\bigg{)}^{\mathbf{1}_{\{s_{t}=s\}}}w^{n}_{t},$ (2)
where $\mathbf{1}_{\{s_{t}=s\}}$ is the indicator function taking on the value $1$ if the state realized at time $t$ is state $s$, i.e., $s_{t}=s$, and equals zero otherwise. Equivalently, the wealth at time $t>0$ is
$w^{n}_{t}=\prod_{\tau=1}^{t}\prod_{s=1}^{S}\bigg{(}\frac{\alpha^{n}_{s\tau}}{p_{s\tau}}\bigg{)}^{\mathbf{1}_{\{s_{\tau}=s\}}}w^{n}_{0}.$ (3)
Directly analyzing individual wealth processes is potentially complex as they depend on prices which are endogenous. Instead, we analyze relative wealths. Let $r^{nm}_{t}$ be the ratio of the wealth of agent $n$ to the wealth of agent $m$ at time $t$. Prices cancel, so wealth ratios follow
$r^{nm}_{t+1}$ $=\prod_{s=1}^{S}\bigg{(}\frac{\alpha^{n}_{st}}{\alpha^{m}_{st}}\bigg{)}^{\mathbf{1}_{\{s_{t}=s\}}}r^{nm}_{t},$ (4)
$r^{nm}_{T}$ $=\prod_{t=1}^{T}\prod_{s=1}^{S}\bigg{(}\frac{\alpha^{n}_{st}}{\alpha^{m}_{st}}\bigg{)}^{\mathbf{1}_{\{s_{t}=s\}}}r^{nm}_{0}.$ (5)
Taking the logarithm of both sides yields
$\log(r^{nm}_{T})=\sum_{t=1}^{T}\sum_{s=1}^{S}\mathbf{1}_{\{s_{t}=s\}}\log(\frac{\alpha^{n}_{st}}{\alpha^{m}_{st}})+\log(r^{nm}_{0}).$ (6)
To provide some insight into how wealth ratios evolve, it is useful to first consider a special case in which investment sequences are constant over time: $\alpha^{n}_{t}=\alpha^{n}$ for all $t$. With constant investments, dividing both sides of equation (6) by $T$, taking the limit as $T\to\infty$, omitting the constant term $\log(r_{0}^{nm})$ that vanishes in the limit, and applying the Law of Large Numbers yields [^4] Note that if $\alpha^{n}_{st}$ were a randomized strategy with expectation $\alpha^{n}$ we could not simply take the expectation, but by Jensen’s inequality, we would get a wealth ratio bounded by the entropy difference $\lim_{T\to\infty}T^{-1}\log(r^{nm}_{T})\leq I_{q}(\alpha^{m})-I_{q}(\alpha^{n})$.
$\lim_{T\to\infty}T^{-1}\log(r^{nm}_{T})=\sum_{s=1}^{S}q_{s}\log(\frac{\alpha^{n}_{s}}{\alpha^{m}_{s}}).$ (7)
To interpret equation (7) it is useful to define the relative entropy (a.k.a. KL-divergence) of investment rule $\alpha$ relative to the probability distribution $q$ as
$I_{q}(\alpha)=\sum_{s=1}^{S}q_{s}\log(\frac{q_{s}}{\alpha_{s}}).$ (8)
Then, rewriting equation (7) using relative entropies we have
$\lim_{T\to\infty}T^{-1}\log(r^{nm}_{T})=I_{q}(\alpha^{m})-I_{q}(\alpha^{n}).$ (9)
Thus, in the constant investment rules environment, the fate of agent $n$ depends on $I_{q}(\alpha^{n})$ and $I_{q}(\alpha^{m})$ for all other traders $m$. For example, if $n$ has the unique lowest relative entropy then $n$’s wealth converges to $1$ and all other agents’ wealth converge to 00. Entropy is minimized at $\alpha=q$, so if a single agent invests according to $q$, that agent’s wealth share converges to $1$.
In the next two sections, we consider Bayesian learners and regret minimizers. Their investment rules are not constant, so we need to generalize this analysis to allow for non-constant rules.
Before considering specific learning strategies, it is useful to note that due to the multiplicative nature of wealth evolution (see Equation (2)), any learning strategy that assigns zero weight to any state risks losing all wealth, with no possibility of recovery; formally:
Definition 2.1.
An agent $n$ survives in the market if there exists a constant market share $c$ such that
$\lim_{T\rightarrow\infty}\Pr\left[\operatorname{\mathbb{E}}[w^{n}_{T}]>c\right]=1.$
We say that agent $n$ vanishes if $n$’s expected wealth converges to zero in the limit. That is,
$\lim_{T\rightarrow\infty}\operatorname{\mathbb{E}}[w^{n}_{T}]=0\text{ with probability }1.$
Observation 2.2.
Let $\{\alpha^{n}_{t}\}_{t=0}^{\infty}$ be a sequence of investment profiles for agent $n$. If there exists a atate $s\in S$ for which $\alpha^{n}_{st}=0$ infinitely often, then agent $n$ vanishes from the market.
This straightforward observation immediately rules out certain learning approaches, such as those that treat asset selection as pure actions. It demonstrates that any reasonable learning strategy must be restricted to investment profiles with full support on the assets. Specifically, all learning strategies we consider will assign at least a small, constant probability $\Delta$ to each state.
3 Comparing Learners
As we noted above, an investor who knows the probabilities $q_{s}$ maximizes the expected growth rate of their wealth at each time $t$ using the investment rule $\alpha^{n}_{t}=q$. Next, consider an agent $n$ with a sequence of investment profiles $\alpha_{t}^{n}$. We will first compare the wealth of this agent to that of an agent who uses the investment rule $\alpha=q$, and then, we will compare different types of learning agents.
As we saw in the previous section in Equation (6), the log of the wealth ratio of an agent at time $T$ compared to an investor using the optimal $\alpha=q$ rule can be expressed as
$\sum_{t=1}^{T}\sum_{s=1}^{S}\mathbf{1}_{\{s_{t}=s\}}\log\alpha^{n}_{st}-\sum_{t=1}^{T}\sum_{s=1}^{S}\mathbf{1}_{\{s_{t}=s\}}\log q_{s}+\log(r^{nm}_{0}).$ (10)
The main goal of this paper is to compare different theories on how to best learn to invest. One theory is Bayesian learning, assuming that the investor $n$ has a Bayesian prior over the possible asset return probabilities. At each step, a Bayesian learner uses the expected return probability implied by their beliefs, updating their beliefs according to the observed states by using Bayes rule. One can also use multi-armed bandit learning methods given such a prior of possible asset return distributions, and have the agent invest using one of the distributions in its prior at each step, aiming to learn which one is the best. A different theory would consider the maximization problem of Equation (10) as a concave maximization problem and use no-regret learning algorithms to optimize the value.
From the latter two perspectives, if the difference between the two sums in Equation (10) is negative, this is a form of regret [^5] This is regret only with respect to $q$; while $q$ is optimal in expectation, it may not be the best fixed strategy in hindsight for a given sequence of state realizations. We extend on this point later; see Theorem 5.2. against a fixed strategy $\alpha=q$. First, consider an investor with a fixed number of possible models for $q$, $\{\theta^{1},\ldots,\theta^{K}\}$. This formulation of the problem would allow us to consider an investor using a classical multi-arm bandit approach, such as UCB (see, e.g., [75]): the investor can invest using one of the $\theta^{k}$ as investment rules, and use no-regret methods to update their belief on the expected value of each of the options. Assuming the correct distribution $q$ is among the options considered, we can express the gap in value between any one choice $\theta^{k}$ and the true distribution using Equation (9) as
$I_{q}(\theta^{k})-I_{q}(q)=I_{q}(\theta^{k}).$ (11)
Classical learning algorithms, such as UCB, then guarantee regret against the optimal strategy $q$ depending inversely on the above gap that is at most
$O(\log T)\cdot\sum_{k:\theta^{k}\neq q}\frac{1}{I_{q}(\theta^{k})}.$ (12)
In the next section, we consider a Bayesian investor who not only has a finite list of possible $q$ vectors, but also has a prior over these models, and uses Bayes rule to update their prior given the outcomes seen in previous rounds.
Alternatively, we can use learning without a list of options or a prior, and hence no longer depend on the assumption that the correct probability vector is one of the options considered. Let us define the utility of an investor at time $t$ with investment strategy $\alpha$ as
$U^{t}(\alpha)=\sum_{s}\mathbf{1}_{\{s_{t}=s\}}\log\alpha_{st}.$ (13)
$U^{t}(\alpha)$ is the rate of change in the investor’s relative wealth in this period compared to the investor using $\alpha=q$. Notice that this is a concave function of $\alpha$ for any outcome of the random variables. For an investor $n$ with investment sequence $\alpha^{n}_{1:T}=(\alpha^{n}_{1},\ldots,\alpha^{n}_{T})$, the value the investor is aiming to optimize is
$\sum_{t=1}^{T}U^{t}(\alpha^{n}_{t}).$ (14)
Using these definitions, the classical regret from no-regret learning for an investment sequence $\alpha_{1:T}^{n}$ at time $T$ is defined as
$R^{T}(\alpha_{1:T}^{n})=max_{\alpha\in\Delta_{S}}\sum_{t=1}^{T}U^{t}(\alpha)-\sum_{t=1}^{T}U^{t}(\alpha^{n}_{t}).$ (15)
Note that the regret of $\alpha^{n}_{1:T}$ is at least as large as its regret against the investor with the optimal fixed rule $\alpha=q$. The maximum fixed rule $\alpha$ used in the definition is the best fixed investment rule with hindsight, having seen the outcome of all choices and random state realizations, which is not an implementable investment rule even for an investor who knows the probabilities. So even the investor with knowledge of the probabilities and using $\alpha^{n}_{t}=q$ for all $t$ will have some regret based on the outcome of random events (see Section 5 and Theorem 5.2).
Next, we reconsider two investors, $n$ and $m$, with investment sequences $\alpha_{1:T}^{n}$ and $\alpha_{1:T}^{m}$, respectively. By Equation (6), we can express their time-$T$ wealth-shares using regret notation as
$\log(r^{nm}_{T})=\sum_{t=1}^{T}U^{t}(\alpha^{n}_{t})-\sum_{t=1}^{T}U^{t}(\alpha^{m}_{t})+\log(r^{nm}_{0})=R^{T}(\alpha_{1:T}^{m})-R^{T}(\alpha_{1:T}^{n})+\log(r^{nm}_{0}).$ (16)
We note that despite this very strong benchmark for regret, and without any prior, the convex optimization literature offers very strong guarantees on expected regret. For any concave function $U^{t}(\alpha)$, a simple gradient-decent based algorithm guarantees regret at most $\sqrt{T}$ over $T$ periods, while for the $\log$ based utility function (that is Exp-Convex) considered here, second order optimization methods guarantee a $O(\log T)$ regret bound [15] (see also, [26] Theorem 3.3 or [56] Theorem 4.4.)
The central question we ask in this paper is: What learning algorithms are better in ensuring that the player’s wealth remains high, or at the very least, does not vanish? How do Bayesian learners and convex optimizers compare in terms of relative wealth in this game? Using equation (16) we get the following result, connecting regret and survival in the market.
Theorem 3.1.
Suppose that there are two agents, $n$ and $m$, in the market with regrets $R^{n}(T)$ and $R^{m}(T)$ respectively. If $\ \lim_{T\rightarrow\infty}R^{n}(T)-R^{m}(T)=\infty$, then agent $n$ vanishes from the market.
Agent $n$ survives [^6] Note that survival and vanishing are probability $1$ statements. Thus, survival is not simply the negation of vanishing. if and only if for every agent $m\neq n$ it holds that $lim_{T\rightarrow\infty}R^{n}(T)-R^{m}(T)<\infty$.
Proof.
This result follows from the above analysis, leading to Equation (16). For the first part, we have $R^{n}(T)-R^{m}(T)=-\log(r^{nm})-\log(r_{0}^{nm})=\log(r^{mn})+const$. If the left hand side diverges to +$\infty$+ ∞, then $r^{mn}=w^{m}/w^{n}$ diverges as well, which is only possible if $w^{n}\rightarrow 0$. For the second part, clearly this is a necessary condition due to the first part of the theorem: if an agent vanishes, it does not survive. To see that this condition is sufficient, if there exists a constant $c>0$ such that $R^{n}(T)-R^{m}(T)<c$ at all times for all $m\neq n$, then the wealth ratio $r^{mn}$ with any other player is bounded, and thus $w^{n}$ is bounded from below. ∎
Recall that the entropy of the investment rule using the correct distribution of states is $I_{q}(q)=0$. Also, recall from equation (11) that the gap in expected reward between this and a different investment rule $q^{\prime}$ is the entropy $I_{q}(q^{\prime})$. This shows that the wrong fixed strategy will have linear regret $\Theta(I_{q}(q^{\prime})T)$. Learners with $O(\sqrt{T})$ do better and learners with only $O(\log T)$ regret do even better. Note that the constant in the regret term will matter. The investor with regret higher by only a constant factor will also vanish! But how well do all these learners do against a Bayesian learner?
4 Bayesian Learners
Next, we consider the standard Bayesian learning results. We restate and summarize in this section known results for completeness and to provide a framework for our comparison of Bayesian and no-regret learning.
Consider a Bayesian watching the process on states in the market. Suppose that the Bayesian does not know the probability $q$, and instead considers a finite number of possibilities for $q$: models $\theta^{1},\dots,\theta^{K}$. We assume that $\theta^{k}_{s}\geq\Delta$ for all $k$ and $s$. See the remark at the end of this section for connection to results for Bayesian learners with a continuum prior. The Bayesian has a prior $\lambda_{0}=(\lambda^{1}_{0},\dots,\lambda^{K}_{0})$ with $\lambda^{k}_{0}>0$ for all $k$. Let $n_{t}^{s}$ be the number of times that state $s$ occurs by time $t$. The posterior probability of model $\theta^{k}$ is, by Bayes rule,
$\lambda^{k}_{t}$ $=\frac{\lambda^{k}_{0}\prod_{s}(\theta^{k}_{s})^{n_{t}^{s}}}{\sum_{l}\lambda^{l}_{0}\prod_{s}(\theta^{l}_{s})^{n_{t}^{s}}}.$ (17)
Thus, for any models $k$ and $l$
$\frac{\lambda^{k}_{t}}{\lambda^{l}_{t}}$ $=\frac{\lambda^{k}_{0}\prod_{s}(\theta^{k}_{s})^{n_{t}^{s}}}{\lambda^{l}_{0}\prod_{s}(\theta^{l}_{s})^{n_{t}^{s}}},$ (18)
$\log\frac{\lambda^{k}_{t}}{\lambda^{l}_{t}}$ $=\log\frac{\lambda^{k}_{0}}{\lambda^{l}_{0}}+\sum_{s}n_{t}^{s}\log(\theta^{k}_{s})-\sum_{s}n_{t}^{s}\log(\theta^{l}_{s}).$ (19)
Using the law of large numbers and the definition of relative entropy, we have, almost surely,
$\lim_{t}\frac{1}{t}\log\frac{\lambda^{k}_{t}}{\lambda^{l}_{t}}=I_{q}(\theta^{l})-I_{q}(\theta^{k}).$ (20)
If model $k$ is correct, $\theta^{k}=q$, then $I_{q}(\theta^{k})=0$, so for any other model $\theta^{l}$ with $I_{q}(\theta^{l})\neq 0$ we have $\lambda^{l}_{t}\to 0$ and $\lambda^{k}_{t}\to 1$ almost surely. Equation (20) shows that the log of posterior probabilities converges linearly in time, and thus that the ratio of posterior probabilities converges at an exponential rate.
The Bayesian learns the true model if it is one of the models considered. Alternatively, suppose that no model considered by the Bayesian is correct, but that there is a unique model with minimum entropy relative to the true model, i.e., there is a $k^{*}$ such that $0<I_{q}(\theta^{k^{*}})<I_{q}(\theta^{k})$ for all $k\neq k^{*}$. Then it follows from equation (20) that $\lambda_{t}^{k^{*}}\to 1$ almost surely.
These convergence results can be summarized in the following proposition.
Proposition 4.1.
Let $\Theta$ denote the support of a Bayesian learner’s prior, and let $\theta^{*}=\operatorname*{argmin}_{\theta\in\Theta}I_{q}(\theta)$. The Bayesian’s beliefs converge to $\theta^{*}$ almost surely.
The Bayesian’s time $t$ predicted probability on states is $q_{t}=\sum_{k}\lambda^{k}_{t}\theta^{k}$. If one of the models the Bayesian considers is correct, then this predicted probability converges at an exponential rate to $q$
4.1 An Investment Rule for the Bayesian
We next generate an optimal investment rule for the Bayesian from the predicted probability on states. A Bayesian trader $n$ with predicted probability $q^{n}_{t}$ who wants to maximize the expected growth rate of wealth selects at time $t$ an investment rule that solves:
$Max_{\alpha^{n}_{t}\geq 0}$ $\sum_{s}{q^{n}_{st}}\log\frac{w^{n}_{t+1}}{w^{n}_{t}}$ (21)
$s.t.$ $\sum_{s}\alpha^{n}_{st}=1,$
$Max_{\alpha^{n}_{t}\geq 0}$ $\sum_{s}{q^{n}_{st}}\log\frac{\alpha^{n}_{st}}{p_{st}}$ (22)
$s.t.$ $\sum_{s}\alpha^{n}_{st}=1.$
The solution is clearly $\alpha^{n}_{t}=q^{n}_{t}$. The Bayesian agent should “bet its beliefs.”
Remark: An alternative foundation for this investment rule can be derived from a Bayesian whose objective is to maximize its expectation of discounted sum of log utility of consumption. Suppose the agent begins with some wealth $w^{n}_{0}$, prior $\lambda^{n}_{t}$ on models, and discount factor $10<\delta<$ 1. If there is a complete set of Arrow securities at each time, this agent will optimally invest a fraction $\delta$ of its wealth at each time, deriving utility from consuming the rest, and use the investment rule $q^{n}_{t}$. Assuming that all agents have the same discount factor yields results identical to those for an economy where agents maximize expected growth rate of wealth.
4.2 Wealth Dynamics for the Bayesian
A Bayesian agent that maximizes the expected growth rate of its wealth uses the investment rule $\alpha^{n}_{t}=q^{n}_{t}$ at each time $t$ (i.e., betting its beliefs). We consider the evolution of the wealth of the Bayesian relative to any other trader who uses an investment rule that does not depend on future states, i.e., it’s time $t$ value depends only on state realizations through time $t$. The log ratio of their wealths is, from Section 2
$\log(r^{nm}_{T})$ $=\sum_{t=1}^{T}\sum_{s=1}^{S}\mathbf{1}_{\{s_{t}=s\}}\log(\frac{\alpha^{n}_{st}}{q_{s}})-\sum_{t=1}^{T}\sum_{s=1}^{S}\mathbf{1}_{\{s_{t}=s\}}\log(\frac{\alpha^{m}_{st}}{q_{s}})+\log(r^{nm}_{0}).$ (23)
Note that
$\sum_{t=1}^{T}\operatorname{\mathbb{E}}[\log(\frac{\alpha^{n}_{st}}{q_{s}})]-\operatorname{\mathbb{E}}[\log(\frac{\alpha^{m}_{st}}{q_{s}})]$ $=\sum_{t=1}^{T}I_{q}(\alpha^{m}_{t})-\sum_{t=1}^{T}I_{q}(\alpha^{n}_{t})+\log(r^{nm}_{0}).$ (24)
If one of the models that the Bayesian considers is the correct model, $q$, then $I_{q}(\alpha^{n}_{t})$ converges exponentially to 00 almost surely. So in this case, $\sum_{t=1}^{T}I_{q}(\alpha^{n}_{t})$ is finite almost surely. The relative entropy $I_{q}(\alpha^{m}_{t})$ is at least 00 for any investment rule and is 00 if $\alpha^{m}_{t}=q$ for all $t$. Thus, the sums of relative entropies are at least equal to the negative of the finite sum of relative entropies for the Bayesian.
Subtracting the means from the sums in Equation (20) yields a martingale, and the arguments in [17] imply the following conclusion:
Proposition 4.2.
Let $r^{nm}_{T}$ be the wealth ratio of a Bayesian learner $n$ competing against an agent $m$ playing $\alpha^{m}_{t}=q$ for all $t$. Then $\liminf_{T\rightarrow\infty}\operatorname{\mathbb{E}}[r^{nm}_{T}]>0$ almost surely; i.e., the Bayesian survives.
Note that if the correct model is not in the set of models considered by the Bayesian, then the predicted distribution converges to the model closest in relative entropy to $q$, and the sum of relative entropies diverges. In this case, if some investor, m, uses $\alpha^{m}_{t}=q$ for all $t$, then the Bayesian’s wealth converges almost surely to 00.
Remark: In this paper we consider Bayesian learners with a finite set of models. The Bayesian analysis can be extended to priors with a continuum of models in their support, see [31]. In this case, learning occurs at rate $O((\log T)/T)$, where the constant depends on the dimension of the set of models considered. We observe that this convergence rate yields a regret of $O(\log T)$ at time $T$. This implies, by our Theorems 3.1 and 5.2, that a Bayesian learner with continuum prior does not survive in a market when competing with an investor who knows the true underlying process, or with a Bayesian with a finite support. The result about the effect of the dimension of the support of the Bayesian’s prior is known in the economics literature, see [18], but the approach there does not use a measure of regret.
Thus, such a Bayesian will have regret of the same order or magnitude as a no-regret learner. In fact, there is an interesting connection between Bayesian learning with a convex continuum support and no-regret learning using the exponential weights algorithm. One can think of the classical exponential weights algorithm (as used in Theorem 3.3 of [26] or Theorem 4.4. of [54], when choosing their parameters $\eta$, $\alpha$, respectively, as $\eta=\alpha=1$) as doing Bayesian learning with a uniform prior over the convex set of distributions on states. It is interesting to note that for a $d$-dimensional convex prior, [31] show convergence at the rate of $\frac{d}{2}(\log T)/T$, resulting in regret $\frac{d}{2}\log T$ (though they do not discuss regret). By contrast, no-regret learning results prove a $d\log T$ regret bound. While this second bound is a factor of 2 worse, note that the no-regret result is worst case: it does compare to the best single investment strategy in hindsight, but does not assume a steady process underlying the market. Since the two algorithms follow the same investment strategy, the no-regret results prove that Bayesian learning over a convex continuum prior of dimension $d$ also guarantees a worst-case regret of at most $d\log T$ against the best fixed strategy in hindsight, even if the Bayesian’s assumption of a steady underlying process turns out incorrect.
5 The Regret of Bayesian Learners
In this section, we further analyze the relationship between an agent’s regret to its wealth level and long-term survival in market dynamics, and use this analysis to derive the regret level of Bayesian learners. This relationship depends on the competition. In particular, we saw in Theorem 3.1 that the learners who survive (have a positive expected wealth level in the limit) are only those who maintain a constant regret gap relative to the best competitor in the market.
As mentioned in Section 3, using the state distribution $q$ as the investment rule (if $q$ were known) maximizes the expected growth rate among constant strategies. However, even this strategy incurs some regret, as it generally differs from the best strategy in hindsight. We now ask: what is the regret level of using the true state distribution $q$ as the investment rule? This regret level represents the best attainable expected regret. Thus, by Theorem 3.1, evaluating this regret level would serve as the benchmark for determining which regret levels ensure survival (with high probability, see Definition 2.1) against any competitors who do not have information about future state realizations. To set the ground, let us first consider the following definition:
Definition 5.1.
Fix an arbitrary history of state realizations $s_{1},\dots,s_{T}$ and denote the empirical distribution of this history by $\hat{q}$, where $\hat{q}_{s}=\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}_{\{s_{t}=s\}}$. A “magic agent,” indexed also by $\hat{q}$, is an agent playing the (eventual) empirical distribution in every step. That is, $\alpha^{\hat{q}}_{st}=\hat{q}_{s}$ for all $t\in[T]$.
Note that this agent uses information regarding future realizations of random states, which is not available to real agents in a stochastic environment. Next, we derive the regret level of using the correct distribution of states $q$ as the investment strategy.
Theorem 5.2.
An agent using the state distribution $q$ as the investment strategy for all $t$ has a constant expected regret. The expected regret, denoted $R$, depends only on the number of states $S$.
Proof.
Consider an agent indexed by $q$ using the state distribution $q$ as its investment strategy; i.e, $\alpha^{q}_{t}=q$ for all $t$, and we denote the seqeunce by $q_{1:T}$. Agent $1$ competes with a magic agent. The regret of the agent playing strategy $q$ is denoted $R^{T}(q_{1:T})$, and denote its expectation by $R=\operatorname{\mathbb{E}}[R^{T}(q_{1:T})]$. On one hand, using the fact that $\hat{q}$ has no regret by definition, from Equation 16 we have
$R^{T}(q_{1:T})=\log(r^{\hat{q}q}_{T})-\log(r^{\hat{q}q}_{0}).$
On the other hand, taking an expectation of Equation 6, we have
$\operatorname{\mathbb{E}}\big{[}\log(r^{\hat{q}q}_{T})\big{]}$ $=$ $\operatorname{\mathbb{E}}\big{[}\sum_{s}\sum_{t}\mathbf{1}_{\{s_{t}=s\}}\log(\frac{\alpha^{\hat{q}}_{s}}{q_{s}})\big{]}+\log(r^{\hat{q}q}_{0})$
$=$ $T\operatorname{\mathbb{E}}\big{[}\sum_{s}\frac{n_{s}}{T}\log(\frac{n_{s}/T}{q_{s}})\big{]}+\log(r^{\hat{q}q}_{0})=T\cdot\operatorname{\mathbb{E}}[I_{\hat{q}}(q)]+\log(r^{\hat{q}q}_{0}).$
where $n_{s}$ is the empirical count of state $s$ until time $T$. And thus, we get
$R=T\cdot\operatorname{\mathbb{E}}\big{[}I_{\hat{q}}(q)\big{]}.$
That is, the expected regret equals the relative entropy (a.k.a. KL divergence) between an empirical distribution and the state distribution scaled by $T$. The question of survival of this agent now hinges on how does $I_{\hat{q}}(q)$ depend on time. Here, we use a result from information theory, showing that the expected KL divergence of the empirical distribution decreases as the alphabet size divided by the sample size [^7] We are interested in the expectation. [68] also provide concentration bounds for this result. (see, e.g., Lemma 13.2 in [70]). In our case, $\operatorname{\mathbb{E}}[I_{\hat{q}}(q)]\equiv\operatorname{\mathbb{E}}[D_{KL}(\hat{q}||q)]=\frac{S-1}{2T}+o(\frac{1}{T})$. Thus, the expected regret $R$ is constant, depending only on the number of states. ∎
We can now also state the following result about a learner competing with a magic agent.
Corollary 5.3.
An agent indexed by $q$ using the correct distribution $q$ survives when competing against a magic agent. This is since by Theorem 5.2, the strategy $q$ yields constant expected regret $R$, and so the regret difference compared to $\hat{q}$—which by definition yields zero regret—is constant as well. By Theorem 3.1, this implies that $q$ survives.
Our analyses above and in the preceding sections lead to the following characterization of the regret of Bayesian learners and the survival conditions for other learning agents competing with them in the market.
Theorem 5.4.
The following holds for any market parameters:
  1. 1.
    Bayesian learners with any finite-support prior that includes the correct state distribution have expected regret bounded by a constant at all times.
  2. 2.
    A learning agent survives in competition with Bayesians if and only if it has constant regret.
Proof.
By Theorem 5.2, an agent $n$ using $\alpha^{n}_{t}=q$ for all $t$ incurs constant expected regret. By Proposition 4.2, a Bayesian learner whose support includes the distribution $q$ almost surely survives against such a competitor. Next, Theorem 3.1 shows that if some agent survives against a competitor, then the difference between their regret levels must be bounded by a constant at all times. Therefore, the Bayesian competing against the state distribution $q$ almost surely incurs constant regret. Having established the first point, the second follows directly from Theorem 3.1: Since the Bayesian has constant regret, any competitor who survives must also have constant regret. ∎
Before concluding this section, the following lemma provides an expression for the regret of a player as a function of the entropy relative to strategy $q$, using our notation $R$ from Theorem 5.2 for the expected regret of strategy $q$. This will be useful in the following Section on imperfect Bayesians.
Lemma 5.5.
The expected regret of a constant strategy sequence ($\alpha^{n}_{1:T}$ such that $\alpha^{n}_{t}=\alpha^{n}$ for all $t$) is
$R^{T}(\alpha^{n}_{1:T})=T\cdot I_{q}(\alpha^{n})+R.$
Proof.
This follows from the definition of regret as a utility difference, as discussed in Section 3. Specifically, the regret of strategy sequence $\alpha^{n}_{1:T}$ is $R^{T}(\alpha^{n}_{1:T})=U(\alpha^{\hat{q}}_{1:T})-U(\alpha^{n}_{1:T})$, also, $R=\operatorname{\mathbb{E}}[U(\alpha^{\hat{q}}_{1:T})-U(q_{1:T})]$, and denote the regret of $\alpha^{n}_{1:T}$ with respect to the benchmark $q_{1:T}$ as $R^{\alpha}_{q}=U(q_{1:T})-U(\alpha^{n}_{1:T})$. We have $\operatorname{\mathbb{E}}[R^{T}(\alpha^{n}_{1:T})]=\operatorname{\mathbb{E}}[R^{\alpha}_{q}]+R=\operatorname{\mathbb{E}}\Big{[}T\sum_{s}\frac{n_{s}}{T}\log(\frac{q_{s}}{\alpha^{n}_{s}})\Big{]}+R=T\cdot I_{q}(\alpha^{n})+R$. ∎
6 Imperfect Bayesians
In the preceding sections, we saw that a perfect Bayesian is optimal in the sense that it survives almost surely in the market and drives out any player with regret increasing over time. In this section, we explore a scenario where a Bayesian learner suffers small errors, either in its set of prior distributions or in its update rule. We find that, while perfect Bayesians are optimal, Bayesian learning is also very fragile; for example, it is key that the correct probabilities is one of the model they consider. Specifically, we show that an imperfect Bayesian would eventually vanish from the market, either against the state distribution $q$ or in competition with any player with sub-linear regret. This holds even if the Bayesian’s errors are very small and have zero mean.
6.1 Bayesian Learning with Inaccurate Priors
As mentioned in Section 4, a Bayesian learner with $q$ not within the support of its prior converges to using the best strategy within its prior (i.e., the one closest to $q$ in relative entropy), denoted here by $q^{\prime}$. This convergence property leads to the following result.
Theorem 6.1.
A Bayesian agent with the state distribution $q$ not in its support incurs regret linear in $T$, and vanishes from the market in competition against any no-regret learner.
Note that this contrasts with the case of Theorem 5.4, where a perfect Bayesian dominates the market in competition with any learners with regret increasing in time.
Proof.
Let $n$ be the index of a Bayesian agent with $q$ not within its support, and let $q^{\prime}\neq q$ be the strategy within the support that is closest to $q$ in relative entropy. By Proposition 4.1, the strategy used by the Bayesian agent, $\alpha^{n}_{t}$, converges almost surely to $q^{\prime}$. Hence, intuitively, for sufficiently large $t$, the strategies remain bounded away from $q$, leading to linear regret from that point onward. In more detail, define $I_{0}=\frac{1}{2}I_{q}(q^{\prime})>0$. In every sample path $i$ (i.e., an infinite sequence, indexed by $i$, of state realizations and the ensuing sequence of $\alpha^{n}_{t}$), except on a set of measure zero, there exists a time $T_{i}$ such that for all $t\geq T_{i}$, we have $I_{q}(\alpha^{n}_{t})>I_{0}$. The regret accumulated up to time $T_{i}$ is constant (possibly positive or negative). The expected regret difference relative to strategy $q$ at $t>T_{i}$ is $\operatorname{\mathbb{E}}[U^{t}(q)-U^{t}(\alpha^{n}_{t})]\geq I_{0}$, and so the expected regret relative to $q$ until time $T$ is at least $I_{0}\cdot(T-T_{i})+const$. By Theorem 5.2, the strategy $q$ incurs only constant regret, so the regret of the Bayesian agent converging to $q^{\prime}$ grows linearly in $T$. Now, consider a no-regret learner. Such a learner has sublinear worst-case regret and thus sublinear expected regret. By Theorem 3.1, the imperfect Bayesian agent who has linear regret vanishes. ∎
Remark: The possible strategies that the Bayesian can play span the entire convex hull of the prior. Bayesians with a finite prior can (and do) play convex combinations of the prior during the dynamic, but still, they always converge to the vertices—even when a more profitable strategy lies in the interior. In this sense, this convergence property of Bayesians is both a strength (if they have an accurate prior) and a weakness (if they do not).
Typical Survival Time: A question that arises now is what happens when errors are small. In the following analysis, we consider “$\epsilon$-inaccurate Bayesians” who have in their prior a close approximation of $q$, but still with some small error. Specifically, the total variation distance between the best strategy in the prior and the state distribution $q$ satisfies $TV(q^{\prime},q)=\epsilon$ for some $\epsilon>0$.
On the one hand, we know that such an agent asymptotically fails to survive against a regret-minimizing agent for any error $\epsilon$. On the other hand, when $\epsilon$ is small, the inaccurate Bayesian will initially converge very quickly to a distribution close to the correct state distribution, and may capture a significant share of the market value during the early stages.
Our goal now is to analyze the survival time during which the inaccurate Bayesian still maintains a significant market share, and to understand how this survival time is related to the level of error $\epsilon$. This, of course, also depends on the best competitor and how fast they converge.
To estimate this relation and provide an upper bound on the survival time of an inaccurate Bayesian learner, we consider a player using $q^{\prime}$ with $TV(q^{\prime},q)=\epsilon$ (having already converged to this distribution), competing against a second player playing a dynamic strategy with expected regret level increasing as $f(t)$ that is sub-linear with $t$.
First, we would like to express the first player’s error in terms of entropy (i.e., KL divergence with $q$). We can bound the entropy using Pinsker’s inequality [69] for a lower bound and its inverse for an upper bound, were the latter holds for finite distributions with full support, as in our case. We have (where recall that $\Delta=\min_{s}q_{s}$)
$2\epsilon^{2}\leq I_{q}(q^{\prime})\leq\frac{2}{\Delta}\epsilon^{2}.$
Using Lemma 5.5, this can be translated into the regret of the inaccurate agent playing strategy $q^{\prime}$ where $R$ is the regret of an investor using the true state distribution $q$ (see Theorem 5.2):
$I_{q}(q^{\prime})=\frac{1}{T}\big{(}\mathbb{E}[R^{T}(q^{\prime}_{1:T})]-R\big{)}.$
To maximize the survival time, we pick the distribution $q^{\prime}$ to be the one that has the smallest regret, which is when $I_{q}(q^{\prime})$ is the smallest possible: $I_{q}(q^{\prime})=2\epsilon^{2}$. So we have [^8] If the inaccurate agent uses the worst $q^{\prime}$ with error $\epsilon$, the regret calcualtion is similar, but with $\epsilon$ rescaled by a factor of $1/\sqrt{\Delta}$, using the inverse of Pinsker’s inequality: $2\epsilon^{2}\tau/\Delta+R$. The smallest-regret $q^{\prime}$ is used for the bound. $\mathbb{E}[R^{T}(q^{\prime}_{1:T})]=2\epsilon^{2}T+R$. By Equation (16), the regret difference between two players is equal to the log of their wealth ratios plus a constant. By comparing regret levels between the agents, we obtain a bound on the typical survival time $\tau$ up to which the inaccurate player holds, in expectation, more than half the market value; beyond this point, their expected share is less and continues decaying to zero: [^9] An alternative question one may want to ask is, “How accurate should my prior be if I want to hold a share of the market for $T$ steps against the competitor?” For this, one can invert the equation to get $\epsilon=\sqrt{(f(T)-R)/T}$.
$f(\tau)=2\epsilon^{2}\tau+R.$
Remark: Note that $f(t)$ is the competitor’s actual regret. Regret bounds for known learning algorithms (e.g., bandit algorithms such as UCB [75] or gradient-descent and second-order methods like those described in [56]) are typically worst-case bounds. Estimating the expected regret (or the most likely one) in game dynamics in general, or specifically in our investment scenario, is an interesting open problem.
For the special cases that the competitor has constant regret (e.g., an accurate Bayesian), or the competitor has $\log T$ or $\sqrt{T}$ regret we get the following results.
Observation 6.2.
The expected survival time of a Bayesian learner with $\epsilon$-inaccurate prior against any player with constant regret (e.g., a perfect Bayesian) is $O\big{(}\frac{1}{\epsilon^{2}}\big{)}$.
For the the competitor with logarithmic regret, we get an equation of the form $\tau=c_{1}\cdot\exp(c_{2}\epsilon^{2}\tau)$. For small errors $\epsilon$, by expanding the exponent, we have the following bound.
Observation 6.3.
The expected survival time of a Bayesian learner with $\epsilon$-inaccurate prior against any player with logarithmic regret (e.g., a no-regret convex optimizer) is $O\big{(}\frac{1}{\epsilon^{3}}\big{)}$.
Finally, in competition against a learner with regret $f(T)=\sqrt{T}$ the survival time is longer.
Observation 6.4.
The expected survival time of a Bayesian learner with $\epsilon$-inaccurate prior against any player with $\sqrt{T}$ regret is $O\big{(}\frac{1}{\epsilon^{4}}\big{)}$.
The interpretation of the above analysis depends, of course, on the time scale we are interested in. When $T$ is large, which is our main focus (e.g., when trade occurs at high frequency or if investments are long-term), only the long-term survival matters. In this case, a Bayesian with an inaccurate prior will eventually vanish in competition with any learner who converges to the truth (such as a regret minimizer). However, when considering transient states as well, it is possible that a Bayesian with a prior that is inaccurate but relatively close to the true distribution could hold a significant market share for some time before eventually vanishing. See Section 7 for an example of such a scenario.
6.2 Bayesian Learning with Noisy Updates
Next, we consider a different type of imperfect Bayesian learner that does have $q$ in its prior, but in every step performs slightly inaccurate “trembling hand” updates. Also here, we demonstrate that the optimality of Bayesian learning is very fragile, even when the correct distribution lies in the support. To model this, we define a noisy Bayesian learner as one who at each step, either slightly over-weights the current observation or slightly over-weights its current prior, such that, in expectation, both the data and the prior receive the correct weights in every step (i.e., the errors in weight have zero mean).
For concreteness, consider the following scenario of a learner attempting to learn a distribution of states. Suppose there are two states, $s_{t}=0$ with a fixed probability $q\in(0,1)$, and $s_{t}=1$ otherwise. The learner considers two models: $\theta_{a}=q$, which is the correct model, and $\theta_{b}\neq q$, with $\theta_{b}<1$. The log-likelihood is given by
$L(s_{t})=(1-s_{t})\log\Big{(}\frac{\theta_{a}}{\theta_{b}}\Big{)}+s_{t}\log\Big{(}\frac{1-\theta_{a}}{1-\theta_{b}}\Big{)}.$ (25)
Now suppose that in every step $t$ the Bayesian learner performs $\eta$-noisy updates where it over- or under-weights the data compared to the prior with a small excess weight $\eta_{t}$, where $\eta_{t}$ has zero mean. Specifically, for a parameter $\eta>0$, $\eta_{t}=\eta$ with probability $1/2$ and $\eta_{t}=-\eta$ otherwise. We find that even a tiny error with zero mean can have a significant impact, essentially breaking the learning process.
Theorem 6.5.
For any $\eta>0$, the Bayesian learner with $\eta$-noisy updates does not converge.
Proof.
Let $\epsilon>0$. The log-likelihood ratio under the noisy update rule takes the following form:
$\log\Big{(}\frac{P_{t}(\theta_{a})}{P_{t}(\theta_{b})}\Big{)}$ $=(1+\eta_{t})L(s_{t})+(1-\eta_{t})\log\Big{(}\frac{P_{t-1}(\theta_{a})}{P_{t-1}(\theta_{b})}\Big{)}$
$=(1+\eta_{t})\sum_{\tau=0}^{t-1}L(s_{t-\tau})\prod_{k=0}^{\tau}(1-\eta_{k})+\log\Big{(}\frac{P_{0}(\theta_{a})}{P_{0}(\theta_{b})}\Big{)}\prod_{\tau=0}^{t}(1-\eta_{\tau}),$ (26)
where the empty product equals one, and we define $\eta_{0}=0$. The products can be simplified:
$\prod_{\tau=0}^{t}(1-\eta_{\tau})=(1-\eta)^{n_{+}(t)}(1+\eta)^{n_{-}(t)},$
where $n_{+}(t)$ is a binomial random variable counting the number of times $\eta_{\tau\leq t}=\eta$, and $n_{-}(t)=t-n_{+}(t)$. The last term can be written as
$\log\Big{(}\frac{P_{0}(\theta_{a})}{P_{0}(\theta_{b})}\Big{)}\cdot(1-\eta)^{n_{+}(t)}(1+\eta)^{n_{-}(t)}=\log\Big{(}\frac{P_{0}(\theta_{a})}{P_{0}(\theta_{b})}\Big{)}\cdot\Big{[}\Big{(}\frac{1-\eta}{1+\eta}\Big{)}^{n_{+}(t)}\Big{]}\cdot(1+\eta)^{t}.$
Intuitively, since $\operatorname{\mathbb{E}}[n_{+}]=T/2$, this should be close to $(1-\eta^{2})^{\nicefrac{{t}}{{2}}}$ with high probability as $t\rightarrow\infty$, which converges to zero for $\eta<1$. Formally, we have the following claim.
Claim 1.
Let $\epsilon>0$. $\lim_{t\rightarrow\infty}\Pr\Big{[}\left(\frac{1-\eta}{1+\eta}\right)^{n_{+}(t)}\cdot(1+\eta)^{t}>\epsilon\Big{]}=0.$
Proof.
Let $\delta>0$. We need to show that there exists $T$ such that for all $t>T$, the event
$\Big{(}\frac{1-\eta}{1+\eta}\Big{)}^{n_{+}(t)}\cdot(1+\eta)^{t}>\epsilon$
has probability less than $\delta$. Dividing by $(1+\eta)^{t}$ and taking the logarithm, we get
$n_{+}(t)\cdot\big{(}\log(1-\eta)-\log(1+\eta)\big{)}>\log(\epsilon)-t\log(1+\eta).$
Rearranging, we have
$n_{+}(t)<\frac{t\log(1+\eta)-\log(\epsilon)}{\log(1+\eta)-\log(1-\eta)}.$
Denote $c_{\epsilon}=\frac{\log(\epsilon)}{\log(1+\eta)-\log(1-\eta)}$, so
$n_{+}(t)<\frac{\log(1+\eta)}{\log(1+\eta)-\log(1-\eta)}\cdot t-c_{\epsilon}.$
Next, we observe that the coefficient of $t$ is strictly less than $1/2$. To see this, define
$c=\frac{1}{2}-\frac{\log(1+\eta)}{\log(1+\eta)-\log(1-\eta)}=-\frac{\log(1+\eta)+\log(1-\eta)}{2\big{(}\log(1+\eta)-\log(1-\eta)\big{)}}=\frac{-\log(1-\eta^{2})}{\text{Positive Number}}>0.$
Thus, we obtain that the event we want to bound is $n_{+}(t)<\Big{(}\frac{1}{2}-c\Big{)}t-c_{\epsilon}.$ Using Hoeffding’s inequality, using that $\operatorname{\mathbb{E}}[n_{+}]=\nicefrac{{t}}{{2}}$, we bound the probability of this event:
$\Pr\Big{[}\big{|}n_{+}(t)-\nicefrac{{t}}{{2}}\big{|}>ct+c_{\epsilon}\Big{]}<2e^{-\frac{2(ct+c_{\epsilon})^{2}}{4t\eta^{2}}}<2e^{-\frac{1}{2}ct}.$
Setting $T=\lceil\frac{2}{c}\ln\left(\frac{2}{\delta}\right)\rceil$, we conclude that for all $t>T$, the probability of the event
$(1+\eta)^{t}\cdot\Big{(}\frac{1-\eta}{1+\eta}\Big{)}^{n_{+}(t)}>\epsilon$
is less than $\delta$, as required. Hence, the product approaches zero for large $t$: $\lim_{t\rightarrow\infty}\prod_{\tau=1}^{t}(1-\eta_{\tau})=0$ almost surely. ∎
Remark: It is worth noting here that the almost-sure behavior of the product is very different from its expectation, which equals one:
$\sum_{k=0}^{T}\binom{T}{k}p^{k}(1-\eta)^{k}\cdot(1-p)^{T-k}(1+\eta)^{T-k}=\big{(}p(1-\eta)+(1-p)(1+\eta)\big{)}^{T}=1\ \text{ (for $p=1/2$)}.$
The above claim implies that the second term in Equation (6.2) converges to zero almost surely. In other words, the learner forgets the prior, much like a perfect Bayesian learner (albeit at a somewhat faster rate in this case). For the first term of Equation (6.2)
$(1+\eta_{t})\sum_{\tau=0}^{t-1}L(s_{t-\tau})\prod_{k=0}^{\tau}(1-\eta_{k}),$
a similar argument applies: with high probability the product for large values of $\tau$ contributes negligibly to the sum, but there remains a random contribution from the smaller $\tau$ terms. To see why there is no convergence, observe that the coefficient $(1+\eta_{t})$ oscillates randomly around $1$. More formally, let $Y_{\tau}=\prod_{k=0}^{\tau}(1-\eta_{k})$, noting that $Y_{0}=1$, and let $Z_{t}=\sum_{\tau=0}^{t-1}L(s_{t-\tau})Y_{\tau}$. Expanding the sum, we have: $Z_{t}=L(s_{t})Y_{0}+\dots+L(s_{1})Y_{t-1}$ and for the next time step $Z_{t+1}=L(s_{t})Y_{0}+\dots+L(s_{1})Y_{t}$. Since $Y_{t}=(1-\eta_{t})Y_{t-1}$, this simplifies to $Z_{t+1}=(1-\eta_{t})Z_{t}+L(s_{t+1})$. Clearly, this sequence does not converge, as both $L(s_{t+1})$ and $\eta_{t}$ take independent random values at each step. ∎
7 Simulations
Imperfect Bayesian Learning: Consider the following toy example. There is market with two assets where agent 1 is a Bayesian learner and agent 2 is a UCB learner, starting from equal wealth levels. [^10] We use UCB for simplicity. The main factor here is the different regret rates and not the particular algorithm. The distribution of states is $q=(0.7,0.3)$, i.e., state 1 occurs with probability $0.7$, and state 2 otherwise. The Bayesian learner has a prior consisting of three hypotheses: $(0.8,0.2)$, $(0.9,0.1)$, $(0.3,0.7)$, where the first hypothesis is closest to $q$ in relative entropy, but still has an error. The UCB learner considers a similar set of distributions as its action set: $(0.7,0.3)$, $(0.9,0.1)$, $(0.3,0.7)$, but the first of these being the correct distribution $q$. The example demonstrates the trade-off between short-term gains for the inaccurate Bayesian from converging quickly to a distribution close to the correct one, and the long-term dominance of the learner who converges to the correct distribution.
Figure 1(a) shows the step-by-step wealth shares over a duration of $10^{6}$ steps, as well as their moving averages over a window of $10$,$000$ steps for better visualization. After a brief initial phase, the no-regret learner (agent 2) converges to playing the correct distribution almost all the time, and it can be seen that its wealth approaches $1$.
Figure 1(b) depicts the distribution of wealth shares for the agents at each time step during the second half of the simulation. It can be seen that the regret-minimizing agent 2 who is close to the correct distribution in the long run holds almost all the wealth almost all the time. Figure 1(c) depicts the distribution during the initial $10$,$000$ steps, showing a different picture. The Bayesian (agent 1) still managed to hold half of the wealth about half of this time, inline with the typical survival time estimate in Observation 6.3.
(a) Wealth Dynamics. The green oscillations show the step-by-step wealth of player 2. Solid lines show running window averages over the recent 10,000 steps in each time.
(b) Wealth distribution over time steps in the second half of the $10^{6}$ step simulation. The distribution shown in blue is the wealth of player 1 and in green is of player 2.
(c) Wealth distribution over time steps in the 10,000 initial steps of the dynamic. The distribution shown in blue is the wealth of player 1 and in green is of player 2.
Figure 1: Example of wealth dynamics in competition between an inaccurate Bayesian with an error in its support learner and regret minimizing agent who converges to the correct state distribution $q$.
Dynamics with a Constant Regret Difference: Next, we consider an example of two Bayesian learners, both with priors including the correct distribution of states. By Theorem 5.4, their regret levels can differ only by an additive constant, as both are perfect Bayesians with constant regret. Consequently, these agents are bound to coexist in the market. The example demonstrates that while the quality of the priors—provided they include a correct model—does not affect survival, the initial learning phase, which depends on the prior, does influence the steady-state partition of wealth between the agents. In this example, as in the previous one, there are two states, with the underlying distribution given by $q=(0.7,0.3)$. Agent 1 has a prior that includes the correct distribution and two other models: $(0.8,0.2),(0.7,0.3),(0.6,0.4)$. Agent 2’s prior includes one additional model: $(0.8,0.2),(0.7,0.3),(0.6,0.4),(0.5,0.5)$. Both agents start with a uniform prior over the options they consider. Thus, agent 2 places smaller weight initially on the correct model than does agent 1. Figure 2(a) depicts the wealth dynamics of these learners, starting form equal wealth levels.
Figure 2(b) shows the same scenario in competition between two no-regret learners using UCB. While the worst-case bound of UCB suggests that the learner with the larger action space could suffer higher regret (by a constant factor), we see empirically that under a steady stochastic process, the agents are comparable in the long run, where only in the initial learning phase agent 1 with the smaller action set has an advantage. In comparison to the Bayesian learners in Figure 2(a), no-regret dynamics continue to be randomized and “forget” the gap in wealth from the initial learning phase.
Comparison of Bayesian vs. No-Regret Learners: Finally, in Figure 2(c) we present an example of the wealth dynamics of a Bayesian learner (agent 1) in competition with a UCB no-regret learner (agent 2). The system is similar to the ones described above, with two states and $q=(0.7,0.3)$. Both agents consider the models $(0.8,0.2),(0.7,0.3),(0.6,0.4)$, so for both, the correct model $q$ is in the support. As expected according to Theorem 5.4, the perfect Bayesian who has constant regret drives out of the market the no-regret learner whose regret is increasing in time.
(a) Two Bayesian learners.
(b) Two no-regret learners.
(c) Bayesian vs. no-regret learner.
Figure 2: Wealth dynamics in a two-state two-player market. Figure 2(a) shows the competition between two Bayesians where the agents have action sets of different sizes which include the correct model $q$. Agent $1$ considers three models and agent $2$ considers the same three models and one additional model. Figure 2(b) shows the same scenario for no-regret learners both using UCB. Figure 2(c) shows the competition between a Bayesian (agent 1) and a regret-minimizing UCB learner (agent 2) considering the same set of models.
Acknowledgments
Éva Tardos was supported in part by AFOSR grant FA9550-23-1-0410, AFOSR grant FA9550-231-0068, and ONR MURI grant N000142412742.